Find the QD=120-5P and if P=(25/100)*20


1
Expert's answer
2021-10-17T16:51:47-0400

"P=\frac{25}{100}\times20=5."

"Q_D=120-5P"

"Q_D=120-5(5)"

"Q_D=95."